# Watch Your Steps: Observable and Modular Chains of Thought

### v1

@@ 4.1.2 Why measuring performance of a single step is hard
We next turn to more general ways of measuring single-step performance. Unlike the steps in Table 7, in many cases, no such oracle exists, because the output contains natural-language text, or because the utility of the step depends on downstream problem-solving in some complex way.
In general, single-step prompting need not give the same result as executing a step in the context of a larger problem-solving task. Informally we say that a step is non-modular if it behaves differently with a single-step prompt than in the context of a complete trace.
@@
A simple experiment shows that steps can be highly non-modular. We modified the consistent_sports step in the sports_understanding task by removing the first input argument. Recall that this step compares two sports descriptions, which clearly cannot be done without seeing the first one. One might expect this prompt to be less accurate, but evaluated it in the usual way, accuracy is statistically identical to the standard version. [^8] In fact performance us better for the “broken” version, which makes 3 errors instead of 5 on the 190 test cases., as shown in the first two lines of Table 8. The problem, of course, is that the LLM is not restricted to generate step outputs from the step inputs—in generating a trace, the LLM can attend to any previous text. In short, LLM trace-generation is robust enough to generate traces that get the right answer even after this change.
We next explore forcing the LLM to consider only the designated inputs to this step, by modifying the trace: we extract the consistent_sports step, re-execute it with a single-step prompt, and then finally regenerate the remainder of the trace by using the prompting method of Section 2.3.2. Following the terminology of Lanham et al. (2023), we call this process an intervention. Surprisingly, forcing modularity this way still makes no significant different in accuracy!
Closer inspection shows the reason for this. The forced-modular version of the consistent_sports step does give different outputs—it agrees with the in-context step only 56.8% of the time. But in completing the trace, the model LLM often adds an extra consistent_sports call, which is then executed in context, and gives a correct answer. The intervention thus leads to similar accuracy, but results in a slight increase in the average number of steps.
This suggests detecting non-modularity with statistics in addition to accuracy (e.g., trace length), so as to detect compensations the LLM makes to fix forced-modular interventions.
@@ 4.1.3 Defining and measuring modularity
We write a program trace $t^{j}$ as a series of steps
@@
and ask the LLM to predict the rest of the trace, which will potentially change every subsequent $x$, $f$, and $y$, yielding the new trace
$\tilde{t}=f_{1}(x_{1},y_{1})\ldots f_{i}(x_{i},\tilde{y}_{i})\ldots\tilde{f}_{n}(\tilde{x}_{n},\tilde{y}_{n})\tilde{a}$
Let $P({T}|F_{1}=f_{1},X_{1}=x_{1})$ be the distribution of traces produced from input $f_{1},x_{1}$, and let $P(\tilde{T}|F_{1}=f_{1},X_{1}=x_{1})$ be the distribution of traces after this intervention. Assume that the step of completing a a trace prefix does not change the distribution.Below we call this assumption the stable incremental generation assumption. If stable incremental generation holds, and step $i$ is modular, then $P(\tilde{T}|F_{1},X_{1})=P(T|F_{1},X_{1})$.
Similarly, suppose $S$ is any summary statistic derived from a trace, i.e. $s\sim{}P(S|T)$. For example, $S$ could be the length of the trace or the correctness of the trace. (When $S$ is just a deterministic function of $T$ we also write this as $s=g_{s}(T)$.) If we assume stable incremental generation, then if step $i$ is modular this also implies
@@
Specifically, in the experiments below, we take a sample of $m$ inputs $\mathcal{X}_{1}=\{x_{1}^{1},\ldots,x_{1}^{m}\}$ and compute corresponding samples of the summary statistics $\mathcal{S}=\{g_{s}(t^{1},\ldots,g_{s}(t^{m})\}$ and $\tilde{\mathcal{S}}=\{g_{s}(\tilde{t}^{1},\ldots,g_{s}(\tilde{t}^{m})\}$. We detect changes between $\mathcal{S}$ and $\tilde{\mathcal{S}}$ with a sign test for $S_{\mathrm{corr}}$, and a paired $t$-test for $S_{\mathrm{numSteps}}$. For $S_{\mathrm{abTrace}}$, we test if the Jensen-Shannon divergence between $S$ and $\tilde{S}$ is higher than expected using a permutation test (see Section B.5).
@@ 4.1.4 Which tasks are non-modular?
Let’s briefly review the preceding subsections. In Section 4.1.1 we showed that we can prompt the model to accurately execute individual steps in isolation, using the method of Section 2.3.1.
In Section 4.1.3, we defined and motivated modularity: steps are modular if they depend only on their defined inputs, and not other aspects of the trace in which they are embedded.
In Section 4.1.3, we proposed a way to measure modularity of a step by using forced-modularity interventions. A subtlety is that this intervention is based on two changes: (a) replacing the step output with a forced-modular version, and then (b) re-generating the rest of the trace. Hence we can only measure the impact of distributional changes due to (a) if we are confident that (b) does not also change the distribution, as assumption we call stable incremental generation.
In the top of Table 9, we evaluate the result of the forced-modularity intervention on some [^9] We do not use steps that call other steps in this experiment. of the oracle-checkable steps from Table 7 on the dev data. The columns labeled $p$ are the $p$-values from our permutation test, so a small value indicates that the step changed the distribution of the statistic. The column $\hat{p}$ is the $p$-value after a multiple-test correction. [^10] If you perform many tests without this correction, you can expect some non-significant differences to have low $p$-values by chance. The multiple-test correction adjusts for this. Corrected $p$-values below 0.1 are italicised and corrected $p$-values values below 0.05 are boldfaced: these correspond to steps where some statistic is changed by the forced-modularity intervention.
For steps where corrected $p$-values are low, we also tested the assumption of stable incremental generation with another intervention, where only (b) occurs. We call this intervention split-and-generate: we truncate the trace at step $i$ and then re-generate it. We see that the assumption fails for three of the steps, so we cannot conclude that these step is non-modular; hence the final column is labeled “unclear” for them. There is no evidence against the stable incremental generation assumption “solve negations” step, but weak evidence ($p=0.083$) against it for “rewite expressions” step, so we mark the boolean expressions step (with $\hat{p}=0.014$ and stable incremental generation) as non-modular and the other (with $\hat{p}=0.083$ and stable incremental generation) as “likely” to be non-modular.
@@
The situation is quite different for the algorithmically x=simpler NLP tasks. In Table 10 we summarize the results of repeating this experiment on 25 steps from 12 NLP tasks. (We choose steps which were prone to local errors as being relatively more interesting.) Only 3 of these steps show a significant distributional shift with the forced-modularity intervention, and one of these may be explained by a failure of the stable incremental generation assumption.
We conclude that non-modularity is rare in the NLP tasks. We note again that the tradition statistic for interventions of accuracy is not sensitive enough to detect any significant changes.
@@ 4.1.5 Discussion
This observation, of course, depends in part on the specific mocks that we used in these experiments. One disadvantage of this test is that, while one can identify non-modularity (when a null hypothesis is refuted), you cannot ensure modularity—the null might still be refuted if you examine more data. However, it does seem clear that at least most step executions of a step are modular.
@@ 5.1 Understanding CoT and Measuring Faithfulness
There has been a great deal of work on CoT prompting and reasoning in LLMs, as surveyed here Huang and Chang (2022).
The prior work that is most relevant to this paper is analytic work on understanding why CoT works, and why and when it fails. A number of hypotheses have been proposed for CoT’s success: for instance, it has been argued that the longer outputs of CoT provide additional computational power, which is needed to solve computationally complex tasks Feng et al. (2024). More notably, it has been argued that CoT “unlocks” reasoning abilities inherit in language models Wei et al. (2022), which naturally emerge from natural linguistic phenomena Prystawski et al. (2024). This view is supported by the observation that phrases like “let’s think step by step” do often lead to improved LLM performance on many tasks Kojima et al. (2022).
The metaphor of “unlocking” reasoning is very compelling, and may be why there has been little prior research on the degree to which CoT prompts can be used teach models task-specific reasoning methods, and the degree to which LLM behavior can be controlled by presenting method traces (as we showed is possible using Program Trace Prompting). The results of Sections 3.3 and 3.4 illustrate that CoT prompt can also provide a new information to the LLM, as its steps describe a particular strategy for solving a task, which the LLM can then learn.
@@
Here we build on the work of Lanham et al. (2023), where faithfulness is measured by changing a generated explanation (e.g., introducing errors, paraphrasing, etc) and them measuring the impact on accuracy. We extend this work by considering a particular change (forced modularity of a step); using new statistics to measure the impact of a change; and studying the more precise notions of modularity of steps and error locality. Our work is also influenced by Yee et al. (2024), who explore error recovery in CoT models and its effect on measuring faithfulness.
@@ 5.5 Learning programs from traces
Learning from program traces has been studied for many decades in machine learning (e.g., Smith (1984)), and has gone by various names including programming by example and learning by demonstration Cypher and Halbert (1993); Lieberman (2001); Lau et al. (2003). To our knowledge, there has been little work on this topic in the LLM era, although algorithmically produced programs traces have been used to analyze the noise-sensitivity of CoT reasoning Havrilla and Iyer (2024). Our work differs from Havrilla and Iyer (2024) in that we propose program traces as a way of constructing controllable chain-of-thought prompts for end tasks, rather than a tool for generation of analytic data.
@@ 6 Conclusions
We proposed Program Trace Prompting, a variant of CoT prompting in which few-shot CoT demonstrations are wrapped in a semi-formal syntax which (1) identifies and names steps; (2) defines the input/output behavior of steps; and (3) replaces CoT explanations of in-context examples with chains of these formalized steps on the same example. We show that this approach is broadly applicable, in that comparable accuracies can be obtained by many models on many tasks. It also offers some advantages over prior prompting schemes.
@@ B.1 Avoiding overfitting to the test data
One case (Navigation) had a bug in which intermediate results were not printed in some cases. In two cases (Date Understanding and Temporal Sequences) review of the mock algorithm determined it was quite different from a CoT-suggested method, and the PT prompt was replaced with one following a more similar strategy. (Recall the goal of this paper is exploring new ways to observe and analyze CoT explanations, not to explore performance differences on these specific tasks; hence we want to align the algorithms implied by the PT prompts with those implied by the CoT prompts as much as possible.)
@@ B.2 Comparing CoT and PTP
In the experiment of Table 2, it is possible for models to consistently fail on an example. The most common reasons for this are that the LLM’s “guard rails” are triggered and it refuses to produce output (e.g., some of the Causal Reasoning examples involve questions about bombs and shootings); that that model’s RLHF instruction-training makes it reluctant to follow the prompt (the Salient Translation Errors suffered from this); or that the model’s output was too long [^11] Model output limits are trivial to get around by generating output in stages, but we leave implementation of this for later work. (which is more frequently a problem for PTP, which has bulkier prompts and outputs.) The statistics above are computed for the subset of examples for which neither model was blocked. For Salient Translation, we added one additional sentence121212“Do NOT simply answer the question with a multiple-choice answer, always generate the program trace first.” to the prompt template used in Stage 3 to improve coverage. The number of examples blocked was moderate for every task, never more than 15%.
@@
Table 11: Overall results of Table 2 compared with prior work: results (1) are from Li et al. (2023), (2) are from Suzgun et al. (2022). Prior results are over the full BBH test sets, whereas our results are on a sample of approximately 75%, as described above. The chain-of-code method is not a single prompt, and executes some code with Python, so is not directly comparable.

### v2

@@ 4.1.2 Why measuring performance of a single step is hard
We next turn to more general ways of measuring single-step performance. Unlike the steps in Table 7, in most cases, no oracle for step correctness exists, either because the output contains natural-language text, or because the utility of the step depends on downstream problem-solving in some complex way.
In general, single-step prompting need not give the same result as executing a step in the context of a larger problem-solving task (as demonstrated by the last column of Table 7). Informally we say that a step is non-modular if it behaves differently with a single-step prompt than in the context of a complete trace. The performance of a non-modular step is not easily defined, so we must first address the question of detecting and measuring modularity.
@@
A simple experiment shows that steps can be highly non-modular. We modified the consistent_sports step in the sports_understanding task by removing the first input argument. Recall that this step compares two sports descriptions, which clearly cannot be done without seeing the first one. One might expect this prompt to be less accurate, but evaluated it in the usual way, accuracy is statistically identical to the standard version [^9] In fact performance is better for the “broken” version, which makes 3 errors instead of 5 on the 190 test cases., as shown in the first two lines of Table 8. The problem, of course, is that the LLM is not restricted to generate step outputs from the step inputs—in generating step outputs, the LLM can attend to any previous text.
We next explore forcing the LLM to consider only the designated inputs to this step, by modifying the trace: we extract the consistent_sports step, re-execute it with a single-step prompt, and then finally regenerate the remainder of the trace by using the prompting method of Section 2.3.2. Following the terminology of Lanham et al. (2023), we call this process an intervention. Surprisingly, forcing modularity this way still makes no significant different in accuracy! (See the last line of Table 8.)
Closer inspection shows the reason for this. The forced-modular version of the consistent_sports step does give different outputs—it agrees with the in-context step only 56.8% of the time. But when regenerating the remainder of the trace, the LLM often adds an extra consistent_sports call, which is then executed in context, and gives a correct answer. The intervention thus leads to similar accuracy, but results in a slight increase in the average number of steps.
This suggests detecting non-modularity with statistics in addition to accuracy (e.g., trace length), so as to detect compensations the LLM makes to “fix” forced-modular interventions.
@@ 4.1.3 Defining and measuring modularity
We write the program trace $t^{j}$ as a series of steps
@@
and ask the LLM to predict the rest of the trace. This will potentially change every subsequent $x$, $f$, and $y$, yielding the new trace
$\tilde{t}=\ldots f_{i}(x_{i},\tilde{y}_{i})\tilde{f}_{i+1}(\tilde{x}_{i+1},\tilde{y}_{i+1})\ldots\tilde{f}_{n}(\tilde{x}_{n},\tilde{y}_{n})\tilde{a}$
Let $P({T}|F_{1}=f_{1},X_{1}=x_{1})$ be the distribution of traces $t$ produced from input $f_{1},x_{1}$, and let $P(\tilde{T}|F_{1}=f_{1},X_{1}=x_{1})$ be the distribution of traces $\tilde{t}$ after the forced-modularity intervention. Assume that the step of completing a a trace prefix does not change the distribution; below we call this assumption the stable incremental generation assumption. If stable incremental generation holds, and step $i$ is modular, then $P(\tilde{T}|F_{1},X_{1})=P(T|F_{1},X_{1})$.
Similarly, suppose $S$ is any summary statistic derived from a trace, i.e. $s\sim{}P(S|T)$. For example, $S$ could be the length of the trace or the correctness of the trace. If we assume stable incremental generation, then if step $i$ is modular this also implies
@@
Specifically, in the experiments below, we first take a sample of $m$ inputs $\mathcal{X}_{1}=\{x_{1}^{1},\ldots,x_{1}^{m}\}$. Generally $S$ is just a deterministic function of $T$, so we also write this as $s=g_{s}(T)$, and compute corresponding samples of the summary statistics $\mathcal{S}=\{g_{s}(t^{1},\ldots,g_{s}(t^{m})\}$ and $\tilde{\mathcal{S}}=\{g_{s}(\tilde{t}^{1},\ldots,g_{s}(\tilde{t}^{m})\}$. We detect changes between $\mathcal{S}$ and $\tilde{\mathcal{S}}$ with a sign test for $S_{\mathrm{corr}}$, and a paired $t$-test for $S_{\mathrm{numSteps}}$. For $S_{\mathrm{abTrace}}$, we test if the Jensen-Shannon divergence between $S$ and $\tilde{S}$ is higher than expected using a permutation test (see Section B.5).
@@ 4.1.4 Which tasks are non-modular?
Let’s briefly review the preceding subsections.
In Section 4.1.1 we showed that we can prompt the model to accurately execute individual steps in isolation, using the method of Section 2.3.1.
In Section 4.1.3, we defined and motivated modularity: steps are modular if they depend only on their defined inputs, and not other aspects of the trace in which they are embedded. We then proposed a way to measure modularity of a step by using forced-modularity interventions. A subtlety is that this intervention is based on two changes: (a) replacing the step output with a forced-modular version, and then (b) re-generating the rest of the trace. Hence we can only measure the impact of distributional changes due to (a) if we are confident that (b) does not also change the distribution, as assumption we call stable incremental generation.
In the top of Table 9, we evaluate the result of the forced-modularity intervention on some of the oracle-checkable steps [^10] We do not use steps that call other steps in this experiment. from Table 7 on the dev data. The columns labeled $p$ are the $p$-values from our permutation test, so a small value indicates that the step changed the distribution of the statistic. The column $\hat{p}$ is the $p$-value after a multiple-test correction. [^11] If you perform many tests without this correction, you can expect some non-significant differences to have low $p$-values by chance. The multiple-test correction adjusts for this. Corrected $p$-values below 0.1 are italicised and corrected $p$-values values below 0.05 are boldfaced: these correspond to steps where some statistic is changed by the forced-modularity intervention.
For steps where corrected $p$-values are low, we also tested the assumption of stable incremental generation with another intervention, where only (b) occurs. We call this intervention split-and-generate: we truncate the trace at step $i$ and then re-generate it. We see that the assumption fails for three of the steps, so we cannot conclude that these steps are non-modular; hence the final column is labeled “unclear” for them. There is no evidence against the stable incremental generation assumption for the “solve negations” step, but weak evidence ($p=0.083$) against it for “rewite expressions” step, so we mark the boolean expressions step (with $\hat{p}=0.014$ and stable incremental generation) as non-modular and the other (with $\hat{p}=0.083$ and stable incremental generation) as “likely” to be non-modular.
@@
The situation is quite different for the algorithmically simpler NLP tasks. In Table 10 we summarize the results of repeating this experiment on 25 steps from 12 NLP tasks. (We choose steps which were prone to local errors as being relatively more interesting.) Only 3 of these steps show a significant distributional shift with the forced-modularity intervention, and one of these may be explained by a failure of the stable incremental generation assumption.
We conclude that non-modularity is rare in the NLP tasks. We note again that accuracy, the traditional statistic for interventions, is not sensitive enough to detect any changes here.
@@ 4.1.5 Discussion
This observation, of course, depends in part on the specific mocks that we used in these experiments. Another disadvantage of this test is that, while one can identify non-modularity (when a null hypothesis is refuted), you cannot ensure modularity—the null might still be refuted if you examine more data. However, it does seem clear that at least most step executions of a step are modular.
@@ 5.1 Understanding CoT and Measuring Faithfulness
There has been a great deal of work on CoT prompting and reasoning in LLMs, as surveyed here Huang and Chang (2022). The prior work that is most relevant to this paper is analytic work on understanding why CoT works, and why and when it fails. A number of hypotheses have been proposed for CoT’s success: for instance, it has been argued that the longer outputs of CoT provide additional computational power, which is needed to solve computationally complex tasks Feng et al. (2024). More notably, it has been argued that CoT “unlocks” reasoning abilities that emerge in sufficiently large language models Wei et al. (2022); Prystawski et al. (2024). This view is supported by the observation that phrases like “let’s think step by step” often lead to improved LLM performance on many tasks Kojima et al. (2022).
The metaphor of “unlocking” reasoning is very compelling, and may be why there has been little prior research on the degree to which CoT prompts teach LLMs task-specific reasoning methods, and the degree to which LLM behavior can be controlled by presenting more explicit method traces (as we showed is possible using Program Trace Prompting). The results of Sections 3.3 and 3.4 illustrate that CoT prompts also provide function by providing new information to the LLM, as its steps describe a particular strategy for a task.
@@
Here we build on the work of Lanham et al. (2023), where faithfulness is measured by changing a generated explanation (e.g., introducing errors, paraphrasing, etc) and then measuring the impact on accuracy. We extend this work by considering a particular change (forced modularity of a step); using new statistics to measure the impact of a change; and studying the more precise notions of modularity of steps and error locality. Our work is also influenced by Yee et al. (2024), who explore error recovery in CoT models and its effect on measuring faithfulness.
@@ 5.5 Learning programs from traces
Learning from program traces has been studied for many decades in machine learning (e.g., Smith (1984)), and has gone by various names including programming by example and learning by demonstration Cypher and Halbert (1993); Lieberman (2001); Lau et al. (2003). To our knowledge, there has been little work on this topic in the LLM era, although algorithmically produced programs traces have been used to analyze the noise-sensitivity of CoT reasoning Havrilla and Iyer (2024). Our work differs from Havrilla and Iyer (2024) in that we propose program traces as a way of constructing controllable chain-of-thought prompts for end tasks, rather than as a tool for generation of analytic data.
@@ 6 Conclusions
We propose Program Trace Prompting, a variant of CoT prompting in which few-shot CoT demonstrations are wrapped in a semi-formal syntax which (1) identifies and names steps; (2) defines the input/output behavior of steps; and (3) replaces CoT explanations of in-context examples with chains of these formalized steps on the same example. We show that this approach is broadly applicable, in that comparable accuracies can be obtained by many models on many tasks. It also offers some advantages over prior prompting schemes.
@@ B.1 Avoiding overfitting to the test data
One case (Navigation) had a bug in which intermediate results were not printed in some cases. In two cases (Date Understanding and Temporal Sequences) review of the mock algorithm determined it was quite different from a CoT-suggested method, and the PT prompt was replaced with one following a more similar strategy. Recall the goal of this paper is exploring new ways to observe and analyze CoT explanations, not to explore performance differences on these specific tasks; hence we want to align the algorithms implied by the PT prompts with those implied by the CoT prompts as much as possible.
@@ B.2 Comparing CoT and PTP
In the experiment of Table 2, it is possible for models to consistently fail on an example. The most common reasons for this are that the LLM’s “guard rails” are triggered and it refuses to produce output (e.g., some of the Causal Reasoning examples involve questions about bombs and shootings); that that model’s RLHF instruction-training makes it reluctant to follow the prompt (the Salient Translation Errors suffered from this); or that the model’s output was too long [^12] Model output limits are trivial to get around by generating output in stages, but we leave implementation of this for later work. (which is more frequently a problem for PTP, which has bulkier prompts and outputs.) The statistics above are computed for the subset of examples for which neither model was blocked. For Salient Translation, we added one additional sentence131313“Do NOT simply answer the question with a multiple-choice answer, always generate the program trace first.” to the prompt template used in Stage 3 to improve coverage. The number of examples blocked was moderate for every task, never more than 15%.
@@
Table 11: Overall results of Table 2 compared with prior work: results (1) are from Li et al. (2023), (2) are from Suzgun et al. (2022). Prior results are over the full BBH test sets, whereas our results are on a sample of approximately 75%, as described in the text. The chain-of-code method is not a single prompt, and executes some code with Python, so is not directly comparable.